OpenStudy (anonymous):

Match each equation in the left column with its solution in the right column. Match Term Definition 3(x – 6) = 6 A) x = 8 3(2x – 1) = 15 B) x = 2 6(x + 1) = 12 C) x = 3 2(4x + 3) = 22 D) x = 1 WHAT GOOES TO WHICH

OpenStudy (wach):

Try solving/expanding the equations. Do you know how to find x?

OpenStudy (anonymous):

NOT REALLY

OpenStudy (wach):

I will work out the 2nd one, which should show you how to do the equation. From there, you should be able to solve everything else. Ok? :) 3(2x – 1) = 15 First, multiply 3 by everything in the parentheses. 3 x 2x = 6x , 3 x -1 = -3. So: 6x - 3 = 15 add 3 to isolate x to one side 6x = 18 divide by 6 x = 18/6 simplify x = 3 So, the second one is C.

OpenStudy (wach):

Use the same process to solve the other 3. I'm here if you need any assistance :D

OpenStudy (anonymous):

IS THE FIRST 1 A

OpenStudy (anonymous):

IM CONFUSED

OpenStudy (wach):

The first one is A. So you're correct. What are you confused with?

OpenStudy (anonymous):

THE LAST ONE

OpenStudy (anonymous):

IS IT D

hero (hero):

start inserting given values of x into each equation.